### v2

@@ Abstract
Existing language models (LMs) often exhibit a Western-centric bias and struggle to represent diverse cultural knowledge. Previous attempts to address this rely on synthetic data and express cultural knowledge only in English. In this work, we study whether a small amount of human-written, multilingual cultural preference data can improve LMs across various model families and sizes. We first introduce CARE, a multilingual resource of 24.1k responses with human preferences on 2,580 questions about Chinese and Arab cultures, all carefully annotated by native speakers and offering more balanced coverage. Using CARE, we demonstrate that cultural alignment improves existing LMs beyond generic resources without compromising general capabilities. Moreover, we evaluate the cultural awareness of LMs, native speakers, and retrieved web content when queried in different languages. Our experiment reveals regional disparities among LMs, which may also be reflected in the documentation gap: native speakers often take everyday cultural commonsense and social norms for granted, while non-natives are more likely to seek out and document them actively. CARE is publicly available at https://github.com/Guochry/CARE (we plan to add Japanese data in the near future).
CARE: Aligning Language Models for Regional Cultural Awareness
@@ 1 Introduction
The impressive capabilities of multilingual language models (LMs) lead to their widespread use across various cultural groups Ouyang et al. (2022); Dubey et al. (2024). However, many studies have shown that LMs often lack knowledge of diverse cultures and are biased towards Western culture Naous et al. (2023); AlKhamissi et al. (2024); Naous and Xu (2025). While much focus has been on demonstrating these cultural biases, less effort Li et al. (2024a, b); Kirk et al. (2024) has been made to address them. Existing works have largely been limited to expressing cultural knowledge in only English or generating synthetic data with existing LMs for fine-tuning.
Figure 1: Example LM responses to culture-specific questions in Chinese and Arabic. The base LM (Refer to caption) Llama3.1-8B-Instruct fails to respond appropriately, while its aligned versions on CARE generate better responses for Chinese (Refer to caption) and Arab (Refer to caption) cultures.
In this work, we explore the benefits of learning from multilingual multicultural human preference data, and demonstrate how even a modest dataset of 2.6k high-quality human-written samples can significantly enhance LM’s cultural awareness across various model families and sizes.
Specifically, we first introduce the Cultural Alignment REsource (CARE [Uncaptioned image]), comprising 24.1k responses (one human-written and nine LM-generated per question) with native human preference using a 10-point scale on 2580 multilingual questions focusing on Chinese and Arab cultures (825 and 697 in respective languages; additional 1,058 covering other cultures). Compared with existing datasets Yin et al. (2022); Palta and Rudinger (2023); Shi et al. (2024); Chiu et al. (2024b), our data offers more balanced coverage and authentic viewpoints, especially for topics such as social norms and cultural commonsense, where previous works have limited data coverage. To achieve this, we have native speakers both re-annotate existing data resources, and curate new data on those rarely documented yet deeply embedded aspects of daily life.
Using CARE, we demonstrate that multilingual human cultural preference enhances LMs’ cultural awareness through alignment strategies, such as direct preference optimization (DPO, Rafailov et al., 2024), without compromising general capabilities. This cultural alignment proves beneficial beyond standard human preference datasets, with consistent gains across both chain-of-thought (CoT) and role-play prompting settings. We further analyze post-alignment performance across different LMs and find that for cultural alignment to be effective, base LMs need a foundational level of cultural knowledge. That is, cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs. We then examine the impact of cross-cultural alignment, and notice that aligning models with data from multiple cultures results in greater improvements than using single-culture data. This highlights the benefits of expanding geographical coverage in LM alignment.
Moreover, we evaluate the cultural awareness of large LMs (e.g. GPT-4o) in comparison to native speakers and web content. We observe cultural awareness gaps among LMs. For example, the more Chinese-centric Qwen2.5-72B excels on Chinese cultural entities, literacy, and social norms, while GPT-4o and Mistral-Large match or outperform it on opinion and commonsense about Chinese culture, likely due to differences in their training corpora. To contextualize these gaps, we analyze web content across languages. For questions on social norms and cultural commonsense, where native humans exhibit high confidence, Google search in native languages often return fewer relevant answers than in English. This indicates that such knowledge is often unstated in native contexts while more explicitly articulated in foreign languages Yin et al. (2022).
In summary, our key contributions include:
  • We introduce CARE, a multilingual multicultural human preference dataset focusing on Chinese and Arab cultures (§3).
  • We show that cultural alignment with CARE improves LMs across model families (§4).
  • Motivated by LM performance gaps, we analyze web content across languages and reveal documentation disparities (§5).
Figure 2: Example questions in Arabic and Chinese (with English translations) from the 5 cultural context categories in CARE Refer to caption: Cultural Entities, Cultural Opinion, Social Norms, Cultural Commonsense, and Literacy3.1).
@@ 2 Related Work
The widespread use of LMs has sparked research interest in their relevance to diverse cultures Adilazuarda et al. (2024); Shen et al. (2024); Pawar et al. (2024); Liu et al. (2024). Several studies investigate the cultural alignment of LMs by examining their responses to questions from social surveys that reflect human values and attitudes Haerpfer et al. (2021); Cao et al. (2023). It has been consistently shown that LMs favor answers associated with Western culture AlKhamissi et al. (2024); Abdulhai et al. (2023), even when prompted in different languages Masoud et al. (2023); Wang et al. (2023) or after undergoing preference optimization Ryan et al. (2024). While these survey-based studies assess general beliefs on broad cultural topics such as democracy or family values, they have limited coverage of detailed cultural nuances and phenomena such as local customs and daily practices.
Another line of work has developed culture-specific evaluation resources such as knowledge bases of cultural facts Keleg and Magdy (2023); Yin et al. (2022); Zhou et al. (2024b), entity-centric cultural benchmarks Naous et al. (2024), and user self-reported cultural experiences Shi et al. (2024). Other works have also constructed QA datasets for different cultural aspects such as culinary customs Palta and Rudinger (2023), norms Rao et al. (2024); Zhan et al. (2024), social etiquette Chiu et al. (2024a, b), and more Arora et al. (2024); Mousi et al. (2024). However, these are mainly designed for evaluating LMs, often in multi-choice QA or text-infilling formats, and are not readily suitable for improving LMs via human preference learning, which requires free-text QA and human ratings on model outputs.
Several multilingual instruction-tuning Muennighoff et al. (2022); Singh et al. (2024); Alyafeai et al. (2024) and preference resources Lian et al. (2023); Bartolome et al. (2023); Ahmadian et al. (2024) have been developed to improve multilingual performance. While they include culturally relevant samples, they primarily offer general instructions and universal preferences for safety and helpfulness. Few works Li et al. (2024a, b); Kirk et al. (2024) attempt to tune LMs on culturally relevant instructions, but they rely on synthesizing data with existing models and mainly express knowledge in English (more on this in §3.1). This may exacerbate biases in the pre-training data and miss authentic viewpoints. Unlike previous studies, our work focuses on culturally aligning LMs by learning from native speakers’ cultural preferences.
@@ 3 Constructing CARE [Uncaptioned image]
To study preference optimization of LMs for cultural awareness, we introduce CARE, which consists of 2,580 culture-specific questions and 24.1k responses (one human-written and nine LM-generated per question) with human ratings on a 10-point Likert scale. We begin by identifying and addressing the limitations in existing data resources (§3.1), and then further collect new data for cultural alignment (§3.2).
@@ 3.1 Limitations of Existing Data Resources
We start with the Aya dataset Singh et al. (2024), the largest multilingual instruction-tuning resource that contains human-written questions and answers. Though its natural instances collected from fluent native speakers offer authentic viewpoints, only part of its contents focus on culture, and the data quality varies for different languages. After manual inspection and filtering out generic questions (e.g. “How many hearts does an octopus have?”), we yield 1,324 (out of 4,909) question and answer pairs that are culturally relevant in Chinese but only 457 (out of 14,210) samples in Arabic. To expand the Arabic set, we apply the same filtering process to around 2,000 samples from CIDAR Alyafeai et al. (2024), a human-written Arabic instruction dataset, resulting in 162 relevant samples.
Figure 3: Overall coverage of our CARE benchmark per cultural category. CARE provides 16.6$\times$ more coverage of social norm and cultural commonsense questions compared to two existing instruction tuning datasets and four cultural knowledge bases (§3.1).

### v4

@@ Abstract
Language Models (LMs) are typically tuned with human preferences to produce helpful responses, but the impact of preference tuning on the ability to handle culturally diverse queries remains understudied. In this paper, we systematically analyze how native human cultural preferences can be incorporated into the preference learning process to train more culturally aware LMs. We introduce CARE, a multilingual resource containing 3,490 culturally specific questions and 31.7k responses with native judgments. We demonstrate how a modest amount of high-quality native preferences improves cultural awareness across various LMs, outperforming larger generic preference data. Our analyses reveal that models with stronger initial cultural performance benefit more from alignment, leading to gaps among models developed in different regions with varying access to culturally relevant data. CARE will be made publicly available at https://github.com/Guochry/CARE.
CARE: Assessing the Impact of Multilingual Human Preference Learning on Cultural Awareness
@@
Figure 1: Example LM responses to culture-specific questions in the native languages. The base LM (Refer to caption) Llama3.1-8B fails to respond appropriately, while its aligned versions on CARE generate better responses for Chinese (Refer to caption), Arab (Refer to caption), and Japanese (Refer to caption) cultures.
@@ 1 Introduction
After large-scale pre-training, Language Models (LMs) typically undergo a crucial post-training phase (aka “alignment”), which includes supervised fine-tuning and preference tuning, to improve their ability to follow instructions and align with human preferences Ouyang et al. (2022); Rafailov et al. (2024). However, the effects of alignment on the cultural awareness of multilingual LMs (i.e., how well they understand and generate appropriate responses to diverse, culturally relevant queries) remains understudied, partly because frontier open-weight models with multilingual support (e.g., Llama-3, Qwen-2.5) have become available very recently around mid-2024. Researchers have only examined the off-the-shelf aligned LMs (e.g., Aya, Tulu 2) Naous et al. (2024); Ryan et al. (2024), while others explored instruction fine-tuning with culturally relevant data Li et al. (2024a).
Dataset Culture- Topics Manually- Human Languages Format
Specific Curated Preferences
Cultural Evaluation
     Fork (Palta and Rudinger) food-related customs en multiple-choice
     CulturalBench (Chiu et al.) daily life, social etiquette, wider society en multiple-choice
     GeoMLAMA (Yin et al.) cultural commonsense en, fa, hi, sw, zh multiple-choice
     BLEnD (Myung et al.) food, sports, family, education, holidays, work-life 13 languages multiple-choice, free text
     Include (Romanou et al.) general knowledge, social science, professional certifications, etc. 44 languages multiple-choice
     CAMeL (Naous et al.) beverage, clothing, food, location, religion, sports, etc. ar masked prompts
     CANDLE (Nguyen et al.) geography, religion, occupation, food, clothing, etc. en assertions
     LLM-GLOBE (Karinshak et al.) cultural values zh, en multiple-choice, free text
     CulturalTeaming (Chiu et al.) red-teaming cultural questions en multiple-choice
     SHADES (Mitchell et al.) culture-specific stereotypes 16 languages stereotyped statements
     JMMMU (Onohara et al.) art, heritage, history, business, science, medicine, etc. ja multiple-choice
[…]
     Aya (Singh et al.) [Uncaptioned image] news, stories, recipes, scientific texts, etc. 65 languages free text
     CIDAR (Alyafeai et al.) [Uncaptioned image] technology, translation, poetry, grammar, etc. ar free text
     Cameleval (Qian et al.) information provision, reasoning, creative writing, etc. ar free text
     PRISM (Kirk et al.) cross-cultural controversies en multi-turn conversations
     Palm (Alwajih et al.) history, celebrations, sports, literature, etc. ar free text
     CULTUREINSTRUCT (Pham et al.) art, cuisine, cultural norms, festivals, history, etc. en free text
     CultureBank (Shi et al.) social norms, food, communication, festivals, etc. en free text
     CulturePark (Li et al.) human belief, norm, custom en free text
     CultureLLM (Li et al.) value survey 9 languages multiple-choice
CARE (Ours) [Uncaptioned image] cultural entities, opinion, norms, commonsense, literacy ar, ja, zh free text
Table 1: Comparison of datasets for studying LMs’ cultural awareness. CARE is a multilingual, human-annotated preference dataset specifically grounded in culture. ([Uncaptioned image]) indicates resources that include some cultural considerations (in the form of culture-related questions or by recruiting native annotators), while cultural coverage is not their only or primary focus (see § 3.1). Representative examples from each dataset are provided in Appendix A.4.
In this paper, we systematically analyze the extent to which and under what conditions preference optimization (e.g., DPO, KTO, SimPO) can help LMs align with regional user expectations along five key dimensions: cultural entities, cultural commonsense, social norms, opinions, and literacy (see examples in Figure 1). To enable our study, we create [Uncaptioned image] CARE, a multilingual culture-specific dataset with human preference judgments on 31.7k LLM/Human-written responses to 3,490 questions about Chinese, Arab, and Japanese cultures. Questions are drawn from a variety of instruction tuning datasets Singh et al. (2024); Alyafeai et al. (2024) and evaluation resources Chiu et al. (2024b); Palta and Rudinger (2023) that we manually identified as culturally-relevant. We further collected additional questions to improve the coverage on regionally popular but less digitally documented commonsense knowledge and social norms, with help from native speakers who have overseas experience.
Leveraging CARE, we show that a modest, high-quality set of native preferences can enhance LMs’ cultural awareness across model families and sizes, outperforming generic preference data despite having 3–17$\times$ fewer samples, and generalizing to out-of-domain cultural tasks Myung et al. (2024); Romanou et al. (2025). We then conduct controlled studies on various LMs and uncover several important insights for cultural alignment: (1) Base model strength matters: aligning base models with stronger initial cultural awareness such as Qwen2.5-7B and Gemma2-9B lifts scores across all cultural categories, whereas models with weaker initial performance hardly improve after alignment, consistent with the post-training analyses of Ivison et al. (2023). (2) Data efficiency scales: even 25% of CARE can lead to a 74% improvement, and performance rises with scaling to the full dataset. (3) Cross-culture synergy: mixing samples about local and foreign cultures in preference learning improves overall awareness of local cultures (e.g., Chinese 4.69 $\rightarrow$ 4.87, Arabic 2.98 $\rightarrow$ 3.50).
Beyond these insights, we observe that performance on different cultural dimensions varies across model families. For example, the Chinese-centric Qwen2.5-72B excels on Chinese entities and social norms but lags behind GPT-4o on cultural commonsense. This suggests that disparities in data access may lead to uneven cultural competencies across models. The querying language also matters: models consistently perform better when queried about culture-specific phenomena in the native language (i.e., Arabic, Chinese, Japanese) rather than English, with the gap narrowing for cultural commonsense. A likely explanation is that everyday cultural commonsense is often unspoken in native contexts while more explicitly expressed and asked about in foreign languages Yin et al. (2022). Our analysis with search engines further supports this, highlighting the importance of broad cultural and linguistic coverage in training data.
@@ 2 Related Work
Cultural Evaluation.
The widespread use of LMs has sparked research interest in their relevance to diverse cultures Adilazuarda et al. (2024); Shen et al. (2024); Pawar et al. (2024); Liu et al. (2024). Several studies investigate LMs’ alignment to different cultures by examining their responses to social surveys that reflect human values and attitudes Haerpfer et al. (2021); Cao et al. (2023). It has been consistently shown that LMs favor answers associated with Western culture AlKhamissi et al. (2024); Abdulhai et al. (2023), even when prompted in different languages Masoud et al. (2023); Wang et al. (2023); Rystrøm et al. (2025) or after preference optimization Ryan et al. (2024). Another line of work develops culture-specific evaluation resources such as knowledge bases of cultural facts Keleg and Magdy (2023); Yin et al. (2022); Zhou et al. (2024b), entity-centric cultural benchmarks Naous et al. (2024), and user self-reported cultural experiences Shi et al. (2024). Other works have constructed QA datasets for different cultural aspects, such as culinary customs Palta and Rudinger (2023), norms Rao et al. (2024); Zhan et al. (2024), social etiquette Chiu et al. (2024a, b); Qiu et al. (2025), and more Arora et al. (2024); Mousi et al. (2024). However, they are primarily designed for evaluation, often in a multiple-choice QA format, and are not well-suited for aligning LMs through preference optimization, which ideally relies on free-text QA and human preference data. (see Table 1 and §3.1).
Cultural Fine-tuning.
Several resources Muennighoff et al. (2022); Singh et al. (2024); Alyafeai et al. (2024); Lian et al. (2023); Bartolome et al. (2023); Ahmadian et al. (2024) have been developed to improve LMs’ multilingual performance. While they include culturally relevant samples, they primarily offer general instructions and preferences for safety and helpfulness. A few other works have investigated cultural adaptation of LMs via fine-tuning strategies Li et al. (2024a, b); Shi et al. (2024); Kirk et al. (2024); Choenni et al. (2024); Yuan et al. (2024); Pham et al. (2025); Alwajih et al. (2025); Xu et al. (2025); Yao et al. (2025). Unlike past studies (see Table 1), CARE provides human preferences on culture-specific topics in three languages (not translated from English), enabling a direct study of how multilingual preference optimization improves models’ cultural awareness and by how much.
Multilingual Preference Optimization
Existing work aims to improve the alignment of LMs with human preferences across different languages, either by synthesising preference data Pan (2024); Hwang (2024); 2A2I (2024) or developing translation-based methods She et al. (2024); Yang et al. (2024b). Only a few studies have released native multilingual human preference datasets that are neither translated from English nor rated by AI models. Notable examples include OpenAssistant Köpf et al. (2023) and HelpSteer3-Preference Wang et al. (2025b), both of which focus on general-purpose interactions. We focus on culturally-relevant preferences for an in-depth analysis of multilingual multicultural alignment.
Figure 2: Overall coverage per cultural category. CARE provides 16.6$\times$ more social norm and commonsense questions compared to 2 instruction tuning datasets (Aya and CIDAR) and 4 cultural knowledge bases (CultureBank, CultureBench, GeoMLAMA, FORK) combined.
@@ 3 Constructing CARE [Uncaptioned image]
We introduce CARE, a multilingual human preference dataset that consists of 3,490 culture-specific questions and 31.7k human/LLM responses with human preference ratings.
@@ 3.1 Limitations of Existing Data Resources
In this section, we notice limitations of existing resources and take action to fix them. We start with the Aya dataset Singh et al. (2024), the largest multilingual instruction-tuning resource that contains human-written questions and answers. Though its samples are collected from native speakers, only part of its content focuses on culture, as many examples consist of general questions. After manual inspection and filtering out generic questions (e.g. “How many hearts does an octopus have?”), we yield 1,324 (out of 4,909) and 600 (out of 3,264) question and answer pairs that are culturally relevant in Chinese and Japanese but only 457 (out of 14,210) samples in Arabic. To expand the Arabic set, we apply the same filtering process to around 2,000 samples from CIDAR Alyafeai et al. (2024), a human-written Arabic instruction dataset, resulting in 162 relevant samples.
We also examine four existing cultural knowledge bases, namely, CultureBank Shi et al. (2024), CulturalBench Chiu et al. (2024b), GeoMLAMA Yin et al. (2022), and FORK Palta and Rudinger (2023). Since these datasets are exclusively in English, we manually translate them into the corresponding native language. We observe that these knowledge bases focus on broad regional coverage, but supply limited samples per culture: namely, 60, 61, and 50 samples for Chinese, Arab, and Japanese cultures. Besides, since these resources mainly use a multiple-choice or text-infilling format, we converted all answers to free-text responses, where we added further explanations while reducing stereotypes and overgeneralization.
We manually classify all questions into one of five categories: (1) Cultural Entities, where the question asks about culture-specific entities Naous and Xu (2025), (2) Cultural Opinion, where the question asks about a subjective interpretation for a cultural entity, (3) Social Norms, where the question is about social interactions between individuals Huang and Yang (2023), (4) Cultural Commonsense, where the question is about daily phenomena that locals may take for granted Shen et al. (2024), and (5) Literacy, where the question is about the language, proverbs, or slang Wuraola et al. (2024). After filtering all resources, we found an imbalanced coverage across categories, with a notable lack of data on social norms and commonsense, as shown in Figure 2. To align LMs via preference optimization, we also require human judgments of both appropriate and inappropriate responses, which are not offered by existing resources.
3.2 Multilingual Cultural Preference Data
To address the aforementioned shortcomings, we collect new data in three languages as follows:
Social Norm and Commonsense.
To expand the samples on social norms and cultural commonsense, we ask native Chinese, Arabic, and Japanese speakers who are international college students undergoing culture transfer experiences themselves to curate such samples. To help brainstorm, the speakers are instructed to leverage international as well as regional social media platforms or forums (e.g., Twitter, Reddit, Zhihu, RedNote), and search for posts where users describe their culture shift experiences using search keywords such as “most surprised abroad”, “culture shock”, “first time to”, etc. Taking inspiration from such discussions, the speakers create 190, 196, and 260 samples in Chinese, Arabic, and Japanese. This human-curated data is more authentic than synthetic data, preventing some inaccuracy and overgeneralization (e.g., “In China, it is common to drink soup after the main dish.”) we have observed in the existing datasets.
Culture-specific Human Preference Judgments.
Given these culture-specific questions, gathering human preferences on the cultural relevance of responses is crucial to performing cultural alignment through preference optimization. To obtain these cultural preferences, we present the native annotators with the zero-shot responses of 9 different LMs to each question within CARE. We specifically use the instruct version of recent multilingual LMs of Llama3.3-70B Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024a), Gemma2-27B Team et al. (2024), Mistral-Large, and GPT-4o. We also use their smaller-sized Llama3.1-8B, Qwen2.5-7B, Gemma2-9B, and Mistral-7B.
We instruct annotators to rank the generated responses from the most to least culturally appropriate and assign a rating on a 1–10 scale (1: poor $\rightarrow$ 10: excellent). These ratings reflect how well responses align with the cultural expectations of native speakers and are used to construct preference pairs for cultural preference learning (§4.1).
Culture $\alpha$ $r$ ${\rho}$ $\tau$
Arab 0.86 0.90 0.88 0.71
Chinese 0.84 0.89 0.89 0.75
Japanese 0.92 0.93 0.92 0.77
Table 2: Inter-annotator agreement based on Krippendorff’s $\alpha$, Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$.
Human Annotators and Agreement.
To curate the samples in CARE, we recruited 2 Chinese, 2 Arab, and 5 Japanese native speakers who are familiar with the respective cultures. These annotators manually filtered the aforementioned resources and searched online platforms for additional samples, capturing broader community knowledge and ensuring that CARE reflects cultural experiences beyond our annotator pool. The annotators then rated the responses to the questions in CARE.
To ensure the quality of our preference ratings, we additionally recruited 4 Chinese, 1 Arab, and 4 Japanese annotators, who were not involved in the data curation process, to perform double annotation. Each additional annotator was given 75 randomly sampled questions (15 per cultural category), for which they were asked to provide their ratings of model responses. We then compare the ratings of the additional annotators to our initial set of annotators. As shown in Table 2, we find substantial inter-annotator agreements across various metrics, indicating consistent response preferences for questions within our considered cultural categories (see Appendix A.2 for per-category agreements).
Chinese Arabic Japanese Model Entities Opinion Norms C. sense Literacy Average Entities Opinion Norms C. sense Literacy Average Entities Opinion Norms C. sense Literacy Average [Uncaptioned image] Gemma2-9B Vanilla 4.89 8.46 7.55 6.67 4.72 6.49 4.73 6.56 6.33 5.60 3.28 5.33 4.46 6.50 5.67 6.50 2.33 5.09 Aligned[Uncaptioned image] 5.50 8.53 7.90 7.26 5.17 6.89 5.30 6.82 6.93 6.55 3.57 5.84 4.40 6.60 5.80 6.67 2.93 5.28 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Llama3.1-8B Vanilla 3.14 4.16 4.62 3.93 3.03 3.78 4.08 3.62 2.87 3.07 2.07 3.30 2.20 3.93 2.47 2.57 1.97 2.63 Aligned[Uncaptioned image] 3.86 5.90 5.36 5.56 3.69 4.88 4.33 4.43 3.70 4.50 2.36 3.86 3.20 4.67 2.80 3.20 1.67 3.11 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Qwen2.5-7B Vanilla 6.89 7.86 7.48 6.80 7.37 7.28 4.65 5.84 5.44 4.88 2.84 4.61 2.13 5.67 4.13 4.60 2.37 3.78 Aligned[Uncaptioned image] 7.20 8.76 7.66 6.90 7.53 7.61 4.55 6.40 5.55 5.33 3.35 5.06 2.77 5.73 4.30 5.17 1.90 3.97 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Mistral-7B Vanilla 3.03 3.83 3.93 4.38 2.43 3.53 2.56 2.46 2.03 2.13 1.34 2.11 1.70 3.96 2.40 2.48 1.20 2.34 Aligned[Uncaptioned image] 2.43 3.90 4.53 5.00 2.20 3.61 2.60 3.36 2.46 2.10 1.40 2.38 1.83 3.90 2.43 2.33 1.23 2.35 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Average scores (1: poor $\rightarrow$ 10: excellent) in responding to questions related to Chinese culture in Chinese, Arab culture in Arabic, and Japanese culture in Japanese. Performances are presented for vanilla LMs and LMs after cultural alignment using DPO on CARE[Uncaptioned image]. For each LM, the row labeled “Vanilla” corresponds to the original model (gray plot), and “Aligned” is after cultural preference learning (colored plot).
Besides the 5-class cultural categories, we also label each sample with the Associated Culture for experiments in §4.5: Native (questions about the local culture; i.e., Chinese, Arab, or Japanese), Foreign (questions about other cultures; e.g., US, German, etc.), or General (questions that are not specific to a particular culture). For fine-grained regional cultural knowledge evaluation (Appendix D.2), we also annotate samples for their Geographic Scope (sub-nationwide, nationwide, continent-wide, or worldwide). See Appendix A.1 for the annotation guideline.
Data Split.
For the experiments in §4 and §5, we construct culture-specific test sets of 150 questions each for Chinese, Arab, and Japanese cultures, sampling 30 questions per cultural category. Remaining CARE questions (1,513 Chinese, 644 Arabic, 757 Japanese), along with rated responses, form the training set. This training data also includes questions about other cultures (e.g. US, German, etc.) beyond Chinese, Arab, and Japanese.
4 Aligning LMs for Cultural Awareness
We investigate whether, and to what extent, high-quality culture-specific data can enhance LMs. We perform human preference learning for cultural alignment of medium-sized LMs (7$\sim$9B parameters) using CARE that fit in 8 Nvidia A40 GPUs.
4.1 Experiment Setup
Human Preference Optimization.
We experiment with different preference learning algorithms, including DPO Rafailov et al. (2024), KTO Ethayarajh et al. (2024), and SimPO Meng et al. (2024). These algorithms require preference pairs that contrast two responses for each question: one culturally appropriate and one that is less appropriate. We use the highest- and lowest-scored responses from the human preference data collected in CARE. We find that all three offline preference algorithms perform similarly well, as shown in Table 4; subsequently, we present most of the experiment results in this paper with DPO unless otherwise specified.